# NeRAF: 3D Scene Infused Neural Radiance and Acoustic Fields

### v2

@@ 1 Introduction
In response, we introduce NeRAF, a method that generates both novel views and RIRs at new sensor positions by jointly learning radiance and acoustic fields (Figure 1). NeRAF queries the radiance field to construct a voxel representation of the scene that encodes radiance and density. This grid conditions the acoustic field with appearance and geometric 3D priors, without the need for additional annotations. NeRAF’s cross-modal approach benefits both modalities: it achieves state-of-the-art RIR synthesis while being more data-efficient and enhances novel view generation in complex scenes. We validate our method on Sound Spaces (Chen et al., 2020a; 2022b), a simulated dataset, and on RAF (Chen et al., 2024), a real-world dataset. Designed as a Nerfstudio module, NeRAF offers convenient access to audio-visual generation necessary for immersive experiences. Although the approach is cross-modal, each modality can be generated independently and the model can be trained without co-located audio and visual sensors, providing flexibility for various applications. We release NeRAF’s code to the community.
@@ Results.
Few-shot RIR Synthesis.
4.3 Ablation Study
Grid Impact.
@@ 5 Discussion
A primary limitation of our method is its scene dependency, requiring NeRAF to be trained separately for each scene. Future work should explore advancements in generalizable NeRF to develop a unified, generalizable approach for synthesizing audio-visual scenes. The current approach is also limited to static scenes. Addressing the challenge of learning implicit representations for audio-visual scenes with dynamic sound sources would be a significant advancement.
@@ Appendix A Field parametrization
SoundSpaces dataset lacks the complete acoustic field parameterization described in Equation 5. Microphone rotates only around the up-axis, $\theta$, and the source is omni-directional. Thus, NAcF can be expressed as $(\textbf{X}_{m},\textbf{d}_{m}=\theta,\textbf{X}_{s},t)\rightarrow(\text{RIR}_{\text{ch}1},\text{RIR}_{\text{ch}2})(f,t)$.
@@
RAF dataset contains only monaural RIRs. Consequently, we replace the two heads presented in Figure 4 by a single head. The microphone is omni-directional but the source is directional. We can therefore express NAcF as $(\textbf{X}_{m},\textbf{d}_{s}=\theta,\textbf{X}_{s},t)\rightarrow\text{RIR}_{\text{mono}}(f,t)$.

### v3

@@ 1 Introduction
In response, we introduce NeRAF, a method that generates both novel views and RIRs at new sensor positions by jointly learning radiance and acoustic fields (Figure 1). NeRAF queries the radiance field to construct a voxel representation of the scene that encodes radiance and density. This grid conditions the acoustic field with appearance and geometric 3D priors, without the need for additional annotations. NeRAF’s cross-modal approach benefits both modalities: it achieves state-of-the-art RIR synthesis while being more data-efficient and enhances novel view generation in complex scenes. We validate our method on Sound Spaces (Chen et al., 2020a; 2022b), a simulated dataset, and on RAF (Chen et al., 2024), a real-world dataset. Nerfstudio (Tancik et al., 2023) is an easy-to-use, modular framework for NeRF. We design NeRAF as a new module to offer convenient access to audio-visual generation necessary for immersive experiences. Although the approach is cross-modal, each modality can be generated independently and the model can be trained without co-located audio and visual sensors, providing flexibility for various applications. We release NeRAF’s code to the community.
@@ 3.4 Neural Acoustic Field
The exact field parametrization adapted to the sensors’ characteristics of each dataset is presented in Appendix A.
@@ Results.
Table 3: Multi-scene training. Comparison of NeRAF performance when trained per scene versus on multiple scenes. Results are averaged across office 4, frl apartment 4 and apartment 2.
Training T60 (%) $\downarrow$ C50 (dB) $\downarrow$ EDT (sec) $\downarrow$
Single-scene 2.29 0.37 0.010
Multi-scene 2.24 0.37 0.010
Generalization on Multiple Scenes.
We evaluate NeRAF’s ability to generalize across multiple scenes. To this end, we train the acoustic field across multiple scenes without increasing the model size. Since Nerfacto is scene-dependent, we first train one NeRF-model for each scene. We populate a 128-grid per scene using the grid sampler and trained NeRFs. These grids are fed into the ResNet3D, which is trained jointly with the acoustic field. Batches are drawn from all scenes during training to ensure global convergence. Table 3 compares single-scene to multi-scene training. The results show similar performance, demonstrating NeRAF’s capability to effectively learn the acoustic field across multiple scenes with a single model. In this setup, the visual modality supports the adaptation to different rooms. However, as NeRF and NAcF are not trained jointly, the visual component does not benefit from cross-modal learning. Detailed results for each scene are provided in Appendix O.
@@
Few-shot RIR Synthesis.
4.3 Ablation Study
Grid Impact.
@@ 5 Discussion
While the neural acoustic field can be trained on multiple scene, it is still necessary to train NeRF separately for each scene. Future work should explore advancements in generalizable NeRF to develop a unified, generalizable approach for synthesizing audio-visual scenes. The current approach is also limited to static scenes. Addressing the challenge of learning implicit representations for audio-visual scenes with dynamic sound sources would be a significant advancement.
@@ Appendix A Field parametrization
SoundSpaces dataset lacks the complete acoustic field parameterization described in Equation 4. Microphone rotates only around the up-axis, $\theta$, and the source is omnidirectional. Thus, NAcF can be expressed as $(\textbf{X}_{m},\textbf{d}_{m}=\theta,\textbf{X}_{s},t)\rightarrow(\text{RIR}_{\text{ch}1},\text{RIR}_{\text{ch}2})(f,t)$.
@@
RAF dataset contains only monaural RIRs. Consequently, we replace the two heads presented in Figure 4 by a single head. The microphone is omnidirectional but the source is directional. We can therefore express NAcF as $(\textbf{X}_{m},\textbf{X}_{s},\textbf{d}_{s}=\theta,t)\rightarrow\text{RIR}_{\text{mono}}(f,t)$.
@@ Appendix I Additional Results Visualization
Appendix K Rationale for using 3D priors
Unlike light, sound propagates as a spherical wavefront in 3D from any point of emission. This means that sound waves spread outward omnidirectionally and interact with the entire 3D room’s geometry, bouncing off surfaces, being absorbed, and reflecting throughout the space. Even with a directional source, which emits sound more strongly in certain directions, the propagation remains omnidirectional in space, influenced by environmental interactions rather than constrained by the initial emission direction.
Previous works (Liang et al., 2023a; b) rely on images (2D) to provide scene information to the acoustic field but we argue that conditioning on 3D information better suits the properties of sound propagation. For example, features like reverberation (linked to the room volume and surfaces) depend on complete 3D volume, not just what’s visible in front of the microphone or camera. By leveraging 3D information, acoustic fields can capture the acoustics of the scene more effectively.
In this paper, we propose to use NeRF as a convenient way for providing 3D scene priors —geometry and appearance— directly to the acoustic field without needing explicit 3D annotations, such as meshes or point clouds, which are complex to obtain.
Appendix L Impact of furnitures
We assess the impact of furnitures nearby a microphone on the predicted RIR. We conducted a first study on the Apartment 1 and Apartment 2 scenes from SoundSpaces, which both include furnished and spacious areas. Given a source, comparing performance across these areas did not reveal a clear trend; differences were observed in both directions. To extend the analysis, we averaged performance across all microphones within each area, regardless of the source. We also conducted this study on a real-scene, Furnished Room, from the RAF dataset. Again, results varied by scene, with no consistent pattern (see table A4). Thus, we find no correlation between a microphone’s proximity to furnitures and prediction accuracy.
Table A4: Performance comparison for microphones located near furniture (Furnished) and in spacious areas (Spacious). No correlation is observed between a microphone’s proximity to furniture and prediction accuracy.
Area T60 $\downarrow$ C50 $\downarrow$ EDT $\downarrow$
Apartment 1
Furnished 2.689 0.598 0.0165
Spacious 2.810 0.521 0.0116
Apartment 2
Furnished 2.382 0.493 0.0138
Spacious 3.470 0.566 0.0102
Furnished Room
Furnished 7.630 0.624 0.0171
Spacious 6.721 0.626 0.0156
Appendix M Unknown Sources
To evaluate NeRAF’s ability to generate RIRs at sources pose unseen during training, we create a new split of the RAF dataset with 85% of sources for training and 15% for testing, maintaining a similar data volume per split as in the original split. Indeed, the original RAF’s split includes overlap in sources between training and testing, though not in source-microphone pairs. This aligns with previous works focusing on RIR prediction at new source-microphone pairs instead of unknown sources (see Table 1). On the unseen source split, NeRAF achieves T60 = 8.04, C50 = 0.76, EDT = 0.028 and STFT error = 0.171, suggesting that NeRAF can generate RIR for unseen sources.
Appendix N Separate training of each modality
Joint training improves image generation in large, complex scenes with sparse visual observations, as shown in table 2 and fig. A6. We assess the impact of separate training of each modality. In that case, a NeRF (without NAcF) is fully trained and used to populate a 3D grid. Then, we train NAcF without the NeRF component. For smaller scenes with sufficient observations, such as office 4, there is no notable difference between vision performances for joint and separate training. Audio generation performance remains equivalent for both approaches, which is expected since the acoustic field in both setups leverages a 3D grid representation of the scene. Detailed results can be found in table A5.
Table A5: Comparison of joint training vs. separate training for the audio and visual components. Joint training improves visual performance in complex scenes with limited training data, while audio generation remains comparable in both cases.
Training T60 $\downarrow$ C50 $\downarrow$ EDT $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Office 4
Joint 1.22 0.25 0.007 25.99 0.921 0.081
Separate 1.22 0.26 0.007 25.94 0.916 0.076
Apartment 1
Joint 2.73 0.57 0.015 27.70 0.915 0.105
Separate 2.71 0.57 0.015 27.00 0.902 0.118
Appendix O Multi-scene Training
We provide multi-scene results when evaluated on each scene in table A6. Each grid is encoded to 1,024 features using the ResNet3D.
Table A6: Multi-scene training. Comparison of NeRAF performances when trained on a single scene against multiple scenes.
Training T60 $\downarrow$ C50 $\downarrow$ EDT $\downarrow$
Office 4
Single-scene 1.22 0.25 0.007
Multi-scene 1.37 0.28 0.007
FRL 4
Single-scene 2.76 0.31 0.010
Multi-scene 2.74 0.31 0.010
Apartment 2
Single-scene 2.88 0.54 0.013
Multi-scene 2.62 0.53 0.013
Mean
Single-scene 2.29 0.37 0.010
Multi-scene 2.24 0.37 0.010